OpenStudy (ilovepurple42043):

use the unit circle to find the inverse function value degrees. cos^-1(sqrt2/2)

OpenStudy (anonymous):

You should ask this question in pre calc you might get a better answer

OpenStudy (ilovepurple42043):

is ther a pre calc section?

Nnesha (nnesha):

is there more information ?

Nnesha (nnesha):

 CAST rule given cos value is positive so its either in first or 4th quadrant

OpenStudy (ilovepurple42043):

ther is sorry. 2 sin theta cos theta+cos theta=0 @Nnesha

OpenStudy (alyssa_xo):

as nnesha said, it's either in the 1st or 4th quadrant http://upload.wikimedia.org/wikipedia/commons/thumb/4/4c/Unit_circle_angles_color.svg/1024px-Unit_circle_angles_color.svg.png reference this

OpenStudy (ilovepurple42043):

yay awesome thanks

Nnesha (nnesha):

\[\huge\rm 2sin \theta \cos \theta + \cos \theta = 0\] take out common factor set it equal to 0 solve for cos and sin